OpenStudy (anonymous):

Find the interquartile range of the data. A. 49 B. 35 C. 16 D. 14

OpenStudy (anonymous):



OpenStudy (anonymous):

that is the data

OpenStudy (anonymous):

This should help: http://www.wikihow.com/Find-the-IQR

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

quickly

OpenStudy (anonymous):

i'll give a medal to anyone that can solve this

OpenStudy (texaschic101):

31,33,35,41,43,46,48,49,49,50 Q2 (median) = (43 + 46)/2 = 89/2 = 44.5 Q1 = 35 Q3 = 49 Interquartile range is Q3 - Q1 = 49 - 35 = 14

OpenStudy (jadeishere):

I just told you all of this... Wow I feel needed...

OpenStudy (anonymous):

thanku now can some one tell me how to give medals

OpenStudy (anonymous):

um huh?

OpenStudy (jadeishere):

Lol select best response /).<

OpenStudy (texaschic101):

click " best response " by the person you want to give the medal to

OpenStudy (anonymous):

i need to give a medal to texaschic101

OpenStudy (anonymous):

there thats it

OpenStudy (texaschic101):

:)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

cool thanku but i have more quesrtions on the way stay tuned

OpenStudy (anonymous):

Find the interquartile range of the data. A. 6 B. 7 C. 8 D. 9